# Quantum Causality: Resolving Simpson’s Paradox with $\mathcal{DO}$-Calculus

### v1

@@ Abstract
Distinguishing correlation from causation is a fundamental challenge in machine intelligence, often representing a critical barrier to building robust and trustworthy systems. While Pearl’s $\mathcal{DO}$-calculus provides a rigorous framework for causal inference, a parallel challenge lies in its physical implementation. Here, we apply and experimentally validate a quantum algorithmic framework for performing causal interventions. Our approach maps causal networks onto quantum circuits where probabilistic links are encoded by controlled-rotation gates, and interventions are realized by a structural remodeling of the circuit—a physical analogue to Pearl’s “graph surgery”. We demonstrate the method’s efficacy by resolving Simpson’s Paradox in a 3-qubit model, and show its scalability by quantifying confounding bias in a 10-qubit healthcare simulation. Critically, we provide a proof-of-principle experimental validation on an IonQ Aria quantum computer, successfully reproducing the paradox and its resolution in the presence of real-world noise. This work establishes a practical pathway for quantum causal inference, offering a new computational tool to address deep-rooted challenges in algorithmic fairness and explainable AI (XAI).
Index Terms:
Causal Inference, Quantum Computing, Simpson’s Paradox, Quantum Machine Learning, $\mathcal{DO}$-Calculus, Quantum Algorithms, Explainable AI (XAI), Algorithmic Fairness.
@@ I Introduction
The remarkable success of modern artificial intelligence (AI) is largely built upon the ability of machine learning models to discern complex patterns and correlations from vast datasets. However, a fundamental limitation of these predominantly data-driven approaches is their inherent inability to distinguish mere correlation from true causation [1]. This distinction is not merely academic; in high-stakes domains such as medical diagnostics, financial modeling, and autonomous systems, decisions based on spurious correlations can lead to unsafe, biased, and unreliable outcomes [2]. Consequently, moving beyond pattern recognition to build models that can reason about cause and effect is a central and pressing challenge for developing the next generation of robust, fair, and trustworthy AI systems.
@@
A stark illustration of this challenge is Simpson’s Paradox [3], a statistical phenomenon where a trend observed across an entire population reverses within its constituent subgroups once the data is partitioned. This phenomenon is particularly pernicious in the causal analysis of observational data, for example, when evaluating the efficacy of a medical treatment on a patient’s outcome. In such scenarios, the paradox emerges when a hidden confounding variable—such as patient demographics that act as a common cause of both the treatment decision and the health outcome—is unevenly distributed across the subgroups. A machine learning model trained on such aggregated data without knowledge of the underlying causal structure would learn the spurious, reversed correlation, leading to fundamentally flawed conclusions. This presents a critical challenge for explainable AI (XAI); if a model’s recommendations, such as in clinical decision support, can be directly contradicted by a more granular analysis of the data, the model’s reliability and trustworthiness are fundamentally undermined. Therefore, resolving such paradoxes is not just a statistical exercise but a prerequisite for building AI systems whose reasoning is transparent and whose conclusions can be trusted.
The theoretical foundation for resolving such paradoxes and reasoning about causal effects is provided by Judea Pearl’s $\mathcal{DO}$-calculus [4]. The $\mathcal{DO}$-calculus offers a formal mathematical framework for computing the results of interventions, such as the interventional probability $P(Y|\mathcal{DO}(X))$, from observational data, provided the underlying causal graph is known. This framework serves as the gold standard for causal inference, allowing researchers to systematically distinguish true causal effects from spurious correlations. While this provides a complete logical framework, it remains a fundamentally abstract statistical procedure. A parallel line of inquiry, therefore, is whether this calculus of interventions can be directly realized within the dynamics of a physical system, offering a different paradigm for computation. Quantum computing, which operates on the principles of physical evolution, presents a natural candidate for such a physical realization, motivating the study of how the abstract “graph surgery” of the $\mathcal{DO}$-calculus can be mapped to concrete operations on a quantum circuit.
In prior work, we introduced a framework for quantum causal inference, establishing how quantum circuits can model causal interventions based on Pearl’s DO-calculus [5]. In this paper, we apply and extend this framework to a canonical challenge in causal reasoning: Simpson’s Paradox. We validate our approach through a series of simulations and a proof-of-principle experiment on quantum hardware. First, using a 3-qubit model, we show that our method successfully resolves a classic Simpson’s Paradox scenario in a high-fidelity simulator. We then reproduce this result on an IonQ Aria trapped-ion quantum processor, confirming the viability of the approach in the presence of real-world hardware noise. Furthermore, to demonstrate scalability, we apply our framework to a complex 10-qubit causal network simulating a healthcare scenario, successfully quantifying the confounding bias that simple statistical stratification fails to correct. Taken together, our results establish quantum causal inference as a promising computational paradigm for tackling long-standing challenges in AI fairness and XAI where complex confounding effects are prevalent. We make no claim of quantum advantage; our goal is to validate the framework on Simpson’s paradox using both simulators and quantum hardware. This v1 preprint does not release code or data; we will make all artifacts available in a subsequent arXiv version (v2).
II Methodology
II-A Causal Framework and the DO-calculus
@@
Simpson’s Paradox is a statistical phenomenon wherein a trend that appears in several groups of data disappears or reverses when these groups are combined. Within the framework of Pearl’s structural causal models (SCMs), this paradox is not a mathematical anomaly but a consequence of misinterpreting causal relationships from observational data. Specifically, it arises when a confounding variable influences both the treatment and the outcome, creating a spurious correlation that can obscure the true causal effect.
@@
In the scenario modeled in this work, the paradox is driven by a confounder, Gender (G), which affects both the Treatment (T) administered and the final Outcome (O). This causal structure can be represented by a Directed Acyclic Graph (DAG), as shown in Fig. 1(a). The confounder G has a direct causal influence on both T ($G\rightarrow T$) and O ($G\rightarrow O$), while T also has a direct influence on O ($T\rightarrow O$). The path $G\rightarrow T$ represents the confounding bias, such as a propensity for one gender to receive a particular treatment more often than the other. To resolve the paradox, we must distinguish between two types of conditional probabilities:
@@
II-B Quantum Implementation of Causal Interventions
@@
By composing a series of these controlled-rotation gates according to the arrows in a given DAG, we construct a quantum circuit that generates a final state distribution equivalent to the joint probability distribution defined by the SCMs. This allows us to construct quantum circuits that model the causal systems under study.
To model the observational distribution, we construct a quantum circuit that directly mirrors the structure of the given causal DAG. Every causal arrow in the graph is translated into a corresponding controlled-rotation gate, as described above, and these gates are composed in sequence to build the full circuit. Crucially, this includes the gates representing the confounding path—in our case, the link from the confounder (Gender) to the treatment (Treatment), $G\rightarrow T$. The resulting quantum circuit, which we term the observational circuit, is thus a physical analogue of the complete, confounded causal system. Executing this circuit and measuring the final state of the qubits yields samples from the observational joint probability distribution $P(O,T,G)$, which inherently contains the statistical biases induced by the confounding variable.
@@
(a) Observational circuit modeling the confounded system.
(b) Interventional circuit implementing the $\mathcal{DO}(T=1)$ operation.
[…]
II-C Specific Circuit Models for Simpson’s Paradox
II-C1 3-Qubit Foundational Model
[…]
II-C2 10-Qubit Scalability Model
@@ II-C Specific Circuit Models for Simpson’s Paradox
III Experiments and Results
III-A Experimental Setup
All experiments were designed and executed within a Python 3 environment, utilizing the Qiskit open-source framework (v1.4.3) [6] for quantum circuit construction and execution. Data analysis and visualization were performed using the standard scientific Python libraries, including Numpy [7], Scipy [8], and Matplotlib [9].
@@ III-A Experimental Setup
III-B Result 1: Simpson’s Paradox on a 3-Qubit Model: Simulation and Experimental Validation
@@
TABLE I: Comparison of 3-qubit Simpson’s Paradox results from the ideal simulator and the IonQ Aria QPU (n=3 trials, 1024 shots per circuit). In both the simulation and the QPU experiment, the treatment effect is positive within the male and female subgroups, but the aggregated observational effect becomes negative, demonstrating the paradox. The quantum causal intervention ($\mathcal{DO}$-operation) successfully resolves this paradox in both cases by revealing the true, positive causal effect. The QPU data qualitatively reproduces the simulation results, while its wider confidence intervals highlight the impact of real-world hardware noise on the statistical outcomes.
@@
Figure 3: Comparison of 3-qubit Simpson’s Paradox results from ideal simulation (blue bars) and the IonQ Aria QPU (orange bars). Both the simulation and the QPU experiment demonstrate the core phenomenon: the treatment effect is positive within the male and female subgroups, but reverses to become negative when aggregated (‘Overall (Obs.)’). In both cases, the quantum causal intervention ($\mathcal{DO}$-operation) successfully resolves the paradox by revealing a robustly positive true causal effect (‘Overall (Causal)’). The QPU results qualitatively reproduce the simulation’s trend, validating the method on real hardware, while the larger error bars on the QPU data reflect the impact of physical noise. Error bars represent 95% confidence intervals, calculated over 30 trials for the simulation and 3 successful trials for the QPU experiment.
@@
III-C Result 2: Scalability on a Complex Network (Simulation)

### v2

# arXiv:2509.00744v2[quant-ph] 28 Feb 2026

@@ Abstract
Distinguishing correlation from causation is a central challenge in machine intelligence, and Pearl’s $\mathcal{DO}$-calculus provides a rigorous symbolic framework for reasoning about interventions. A complementary question is whether such intervention logic can be given executable semantics on physical quantum devices. Our approach maps causal networks onto quantum circuits, where nodes are encoded in qubit registers, probabilistic links are implemented by controlled-rotation gates, and interventions are realized by a structural remodeling of the circuit—a physical analogue of Pearl’s “graph surgery” that we term circuit surgery. We show that, for a family of 3-node confounded treatment models (including a Simpson-type reversal), the post-surgery circuits reproduce exactly the interventional distributions prescribed by the corresponding classical $\mathcal{DO}$-calculus. We then demonstrate a proof-of-principle experimental realization on an IonQ Aria trapped-ion processor and a 10-qubit synthetic healthcare model, observing close agreement between hardware estimates and classical baselines under realistic noise. We do not claim quantum speedup; instead, our contribution is to establish a concrete pathway by which causal graphs and Pearl-style interventions can be represented, executed, and empirically tested within the formalism of quantum circuits.
@@ I Introduction
Distinguishing correlation from genuine cause–effect relationships is a foundational problem in statistics and machine learning, and is increasingly important for data-driven AI systems deployed in the real world [1]. This distinction is not merely academic; in high-stakes domains such as medical diagnostics [2], financial modeling [3], and autonomous systems [4], decisions based on spurious correlations can lead to unsafe, biased, and unreliable outcomes. Consequently, moving beyond pattern recognition to build models that can reason about cause and effect is a central challenge for developing robust, fair, and trustworthy AI systems.
@@
A classic illustration of this challenge is Simpson’s Paradox [5, 6], a statistical phenomenon where a trend observed across an entire population reverses within its constituent subgroups once the data is partitioned. This phenomenon is particularly pernicious in the causal analysis of observational data, for example, when evaluating the efficacy of a medical treatment on a patient’s outcome. In such scenarios, the paradox emerges when a hidden confounding variable—such as patient demographics that act as a common cause of both the treatment decision and the health outcome—is unevenly distributed across the subgroups. A machine learning model trained on such aggregated data without knowledge of the underlying causal structure would learn the spurious, reversed correlation, leading to fundamentally flawed conclusions. From a causal-inference perspective, Simpson’s Paradox thus serves as a canonical stress test: any framework that aspires to compute interventional quantities must explain how such reversals arise from confounding and how they are resolved by appropriate adjustment.
The theoretical foundation for analyzing such situations and, more broadly, for reasoning about causal effects is provided by Pearl’s $\mathcal{DO}$-calculus [7]. The $\mathcal{DO}$-calculus offers a formal framework for computing the results of interventions, such as the interventional probability $P(Y\mid\mathcal{DO}(X))$, from observational data when the underlying causal graph is known. This framework has become a reference point in modern causal inference, allowing researchers to systematically distinguish causal effects from spurious associations. In classical causal inference, this calculus underpins a broad toolbox of identification and estimation techniques that are routinely applied to observational data. At the same time, it is typically realized as an abstract, symbolic procedure on probability distributions rather than as a concrete physical process. A natural question, therefore, is whether this calculus of interventions can also be instantiated in the dynamics of a physical system, providing an alternative computational paradigm for causal reasoning.
Quantum computing presents a natural candidate for such a physical realization, owing to a useful structural analogy between structural causal models (SCMs) and quantum circuits: both are probabilistic at the level of observed outcomes, directed, and defined by local interactions. This analogy motivates the question of how the abstract “graph surgery” of the $\mathcal{DO}$-calculus can be mapped to concrete operations on a quantum circuit.
In prior work [8], we introduced a framework for quantum causal inference, establishing how quantum circuits can model causal interventions based on Pearl’s $\mathcal{DO}$-calculus. In this paper, we apply and extend this framework to a canonical challenge in causal reasoning: Simpson’s Paradox. We validate our approach through a series of simulations and a proof-of-principle experiment on quantum hardware. Our contributions are threefold:
  • We compile binary SCMs into block-diagonal quantum circuits, representing nodes as qubit registers, encoding conditional probabilities via controlled-rotation gates, and giving Pearl’s $\mathcal{DO}$-operator executable semantics through a simple circuit-surgery rule.
  • We instantiate this construction on a 3-qubit confounded treatment model that exhibits a Simpson-type reversal, and validate the $\mathcal{DO}$-calculus resolution both in high-fidelity simulation and on an IonQ Aria trapped-ion noisy intermediate-scale quantum (NISQ) processor, observing close agreement with classical baselines under realistic hardware noise.
  • We apply the same methodology to a 10-qubit synthetic healthcare network with multi-level confounding, using the circuit-level implementation of backdoor adjustment to quantify confounding bias beyond what simple statistical stratification can correct.
Throughout, we make no claim of quantum advantage or new causal identification criteria; instead, we offer an experimental validation of classical $\mathcal{DO}$-calculus on current NISQ hardware and a concrete case study of how SCM-style graph surgery can be executed and empirically tested within the formalism of quantum circuits.
The remainder of this paper is organized as follows. Section II reviews related work on quantum causal models, Born-rule generative models, and quantum analogues of Simpson effects. Section III introduces our causal framework, the circuit-surgery implementation of $\mathcal{DO}$-operations, and the specific 3-qubit and 10-qubit models used in this study. Section IV presents simulation and hardware results, comparing ideal and noisy implementations and quantifying confounding bias. Section V discusses the interpretation of these findings, connections to classical methods, and current limitations and future directions. Finally, Section VI concludes the paper.
II Related Work
Classical causal inference provides a mature toolbox for identifying and estimating causal effects from observational data, grounded in structural causal models and the $\mathcal{DO}$-calculus [1, 7]. Constraint-based discovery algorithms such as the Fast Causal Inference (FCI) procedure [9], functional-model approaches like LiNGAM for linear non-Gaussian systems [10], and identification strategies based on covariate adjustment, instrumental variables, and the front-door criterion [1] offer well-established ways to handle confounding and to recover quantities such as $P(Y\mid\mathcal{DO}(X))$ in fully classical settings. Our goal in this work is not to propose an alternative to these methods, but to investigate how the same intervention logic can be compiled into and experimentally validated on quantum circuits.
In what follows, we briefly review three strands: (i) quantum causal-model frameworks (conditional states, process matrices, and quantum combs), (ii) Born-rule-based quantum-circuit generative models (“Born machines”), and (iii) quantum analogues of Simpson effects driven by interference and contextuality. We then clarify how our approach differs by compiling $\mathcal{DO}$-operations into circuits and validating them on NISQ hardware.
II-A Quantum Causal Modeling
A mature line of work generalizes classical causal models to quantum systems by replacing random variables with quantum systems and conditional probabilities with completely positive (CP) maps or “conditional states.” Foundational formalisms include the conditional-state approach of Leifer and Spekkens (quantum Bayesian inference) [11], the quantum-combs framework that represents multi-time processes and interventions as higher-order maps [12], and the process-matrix formalism that allows for indefinite causal order [13]. On top of these primitives, quantum causal models (QCMs) and discovery algorithms have been proposed, aiming to recover causal structure from data and to define quantum analogues of interventions [14, 15, 16].
Conceptually, QCMs separate causal structure (wiring of CP instruments) from dynamics (the CP maps themselves), mirroring the classical split between a directed acyclic graph (DAG) and its structural equations [15]. In definite-order (acyclic) settings, interventions correspond to replacing instruments along a node (quantum $\mathcal{DO}$ operations), whereas the process-matrix line shows that some quantum processes are causally nonseparable and thus lie beyond any single underlying order [13]. Our work is complementary: we adopt the same operator-theoretic primitives (states, effects, and instruments) to give executable semantics to classical $\mathcal{DO}$-operations, but we do not build a full QCM nor engage with indefinite order; instead, we target a concrete causal task (Simpson’s Paradox) and validate interventional predictions on NISQ hardware with direct comparability to classical baselines.
II-B Born-rule Generative Models (Born Machines)
A parallel line of work leverages the Born rule [17] to define generative models, where a parameterized quantum circuit assigns probabilities to classical data via measurement outcomes. Early formulations—often called Quantum Circuit Born Machines (QCBMs)—optimize circuit parameters so that the Born distribution matches a target dataset, using objectives such as maximum mean discrepancy, likelihood surrogates, or adversarial criteria [18, 19]. Subsequent studies explored scaling, expressivity, and potential advantage of Ising-type Born machines for sampling tasks [20], while also identifying trainability challenges in variational circuits (barren plateaus) [21].
Our use of the Born rule is orthogonal to these generative approaches. We do not train a parameterized circuit to fit a dataset; rather, we use the operator calculus to compile classical $\mathcal{DO}$-operations (intervene, condition, and marginalize) into circuit-level primitives (state preparation, controlled gates, and measurements) and to validate that the resulting interventional probabilities resolve Simpson’s Paradox exactly as predicted by SCM. Phase serves only as a controllable degree of freedom to reproduce classical interaction contrasts.
II-C Quantum Analogues of Simpson’s Paradox
Several works analyze “quantum Simpson effects,” where aggregation reversals arise from noncommuting measurements and amplitude interference rather than classical confounding [22, 23]. These studies sharpen the conceptual distinction between interference-driven and confounder-driven reversals. Our focus is different: we operationalize the classical causal resolution by compiling $\mathcal{DO}$-operations into circuits and verifying them on hardware, thereby providing a physically grounded implementation of interventional logic rather than a contextuality-based account.
III Methodology
In this section we specify the Simpson-type SCMs we study, recall the relevant $\mathcal{DO}$-calculus notions, and describe how these ingredients are compiled into quantum circuits via circuit surgery. For completeness, the formal SCM-to-circuit notation underlying this construction is summarized in Appendix B.
III-A Causal Framework and the $\mathcal{DO}$-calculus
@@
Simpson’s Paradox is a statistical phenomenon wherein a trend that appears in several groups of data disappears or reverses when these groups are combined. This effect is particularly common when analyzing observational data to evaluate, for instance, the efficacy of a medical treatment on a patient’s outcome. Within the framework of Pearl’s structural causal models (SCMs), the paradox is understood not as a mathematical anomaly but as a consequence of misinterpreting these causal relationships. Specifically, it often arises when a hidden confounding variable influences both the treatment and the outcome, creating a spurious correlation that can obscure the true causal effect.
@@
In the scenario modeled in this work, the paradox is driven by a confounder, Gender (G), which affects both the Treatment (T) administered and the final Outcome (O). This causal structure can be represented by a DAG, as shown in Fig. 1(a). The confounder G has a direct causal influence on both T ($G\rightarrow T$) and O ($G\rightarrow O$), while T also has a direct influence on O ($T\rightarrow O$). The path $G\rightarrow T$ represents the confounding bias, such as a propensity for one gender to receive a particular treatment more often than the other. To resolve the paradox, we must distinguish between two types of conditional probabilities:
@@
III-B Quantum Implementation of Causal Interventions
@@
By composing these controlled-rotation gates according to a topological order of the DAG, we construct quantum circuits whose measurement statistics reproduce the SCM’s joint probability distribution. To model the observational distribution, for instance, we build an observational circuit that translates every causal arrow in the graph—including confounding paths like $G\rightarrow T$—into a corresponding gate. Crucially, as our entire construction uses only gates that are block-diagonal in the computational basis, it operates within a classical sector of the operator formalism. This ensures the circuit’s semantics are mathematically identical to the classical SCM, a claim formally proven in Appendix C.
To sample from an interventional distribution, such as $P(O,G|\mathcal{DO}(T=t))$, we implement the $\mathcal{DO}$-operator by performing a procedure based on the project-prepare circuit surgery we introduced in our prior work [8], which for brevity we term “circuit surgery”. Conceptually, this process involves identifying and removing all quantum gates from the observational circuit that correspond to the causal arrows pointing into the variable being intervened upon. Our implementation achieves this by directly constructing a new interventional circuit where these gates are omitted.
@@
III-C Specific Circuit Models for Simpson’s Paradox
(a) Observational circuit modeling the confounded system.
(b) Interventional circuit implementing $\mathcal{DO}(T{=}1)$ via circuit surgery.
Figure 2: Quantum circuits for the 3-qubit Simpson model. (a) The observational circuit encodes all causal links. The dashed box highlights the gates implementing the confounding path ($G\!\to\!T$) targeted for removal via circuit surgery. (b) The resulting interventional circuit, where the confounding path has been removed and $T$ is deterministically prepared to sample from $P(O,G\mid\mathcal{DO}(T{=}1))$. All operations act within the block-diagonal (classical) sector, ensuring Born probabilities coincide with classical SCM semantics.
III-C1 3-Qubit Foundational Model
@@ III-C Specific Circuit Models for Simpson’s Paradox
III-C2 10-Qubit Scalability Model
@@
III-C3 Circuit-Based Resolution of Simpson’s Paradox
To resolve Simpson’s Paradox, one must compute the true causal effect using the backdoor adjustment formula [1], which isolates the effect of $X$ on $Y$ by accounting for the confounder $Z$:
$P(y\mid\mathcal{DO}(x))=\sum_{z}P(y\mid x,z)\,P(z).$
Our quantum framework is designed to be a direct physical realization of this adjustment. The observational circuit first provides the necessary population statistics, such as the distribution $P(z)$. Then, the interventional circuit, created via circuit surgery on $X$, produces samples from the post-interventional distributions, which allows for the calculation of the conditional probabilities $P(y\mid x,z)$.
By combining these results according to the formula, our method is mathematically guaranteed to yield the true causal effect, $P(y\mid\mathcal{DO}(x))$, thereby resolving the paradox. A formal statement and proof that our compiled circuit primitives correctly implement this compositional causal query is provided in Appendix E.
IV Experiments and Results
In this section we report the behavior of our quantum causal intervention framework on both the 3-qubit Simpson-type model and the 10-qubit healthcare network, combining noiseless simulations with execution on the IonQ Aria NISQ processor.
IV-A Experimental Setup
All experiments were designed and executed within a Python 3 environment, utilizing the Qiskit open-source framework (v1.4.3) [24] for quantum circuit construction and execution. Data analysis and visualization were performed using the standard scientific Python libraries, including Numpy [25], Scipy [26], and Matplotlib [27]. The data that support the findings of this study, as well as the source code for the experiments and analysis, are publicly available in a GitHub repository at https://github.com/pilsungk/quantum-simpson.
@@ IV-A Experimental Setup
IV-B Result 1: Simpson’s Paradox on a 3-Qubit Model: Simulation and Experimental Validation
@@
TABLE I: Simpson-type treatment effects in the 3-qubit model from an ideal simulator and the IonQ Aria QPU (3 trials, 1024 shots per circuit). In both backends the treatment effect is positive within each gender subgroup but negative in the aggregated observational analysis, while the quantum $\mathcal{DO}$-intervention recovers a positive causal effect with wider confidence intervals on the QPU reflecting hardware noise.
@@
Figure 3: Comparison of 3-qubit Simpson’s Paradox results from ideal simulation (blue bars) and the IonQ Aria QPU (orange bars). Both the simulation and the QPU experiment demonstrate the core phenomenon: the treatment effect is positive within the male and female subgroups, but reverses to become negative when aggregated (‘Overall (Obs.)’). In both cases, the quantum causal intervention ($\mathcal{DO}$-operation) successfully resolves the paradox by revealing a robustly positive true causal effect (‘Overall (Causal)’). Error bars represent 95% confidence intervals, calculated over 30 trials for the simulation and 3 successful trials for the QPU experiment. The QPU results qualitatively reproduce the simulation’s trend, validating the method on real hardware, while the larger error bars on the QPU data reflect the impact of physical noise.
@@
IV-C Result 2: Scalability on a Complex Network (Simulation)